OpenStudy (anonymous):

Square root of 8y^3z^2

OpenStudy (anonymous):

2yz@2y

OpenStudy (anonymous):

or 2yz√2y

OpenStudy (anonymous):

\[\sqrt{8y ^{3}z ^{2}} = \sqrt{2*4*y*y*y*z*z} = 2yz \sqrt{2y}\]

OpenStudy (anonymous):

\[\sqrt[5]{32x ^{8}y ^{3}z ^{2}5}\]

OpenStudy (anonymous):

can some one give me step by spet instructions or explain how to solve this i honestly have no idea